The expected number of typographical errors on a page of a certain magazine is H 0.30 What is the probability that the next page you · read contains: S a) exactly one typographical error? b) at least one typographical error?
Answer #1

Let X denotes the number of typographical errors on a randomly selected page.

X ~ Poisson(mu = 0.75)

The probability mass function of X is

P(X=x) = rac{e^{-0.75}*0.75^x}{x!},x=0,1,2,....

a) The probability that there is exactly one typographical error

P(X = 1)

e-0.15 * 0.75 i

-0.354275

b) The probability that there is at least one typographical error

=P(Xgeq 1)

= 1 - P(X = 0)

1-0.472367

= 0.527633